OpenStudy (anonymous):

A building lot in a city is shaped as a 30-60-90 triangle. The side opposite the 30angle measures 41 feet. a. Find the length of the side of the lot opposite the 60 angle. b. Find the length of the hypotenuse of the triangular lot. c. Find the sine, cosine, and tangent of the 30 angle in the lot. Write your answers as decimals rounded to four decimal places.

OpenStudy (anonymous):

Help please! I'll become a fan and give you best answer whoever answers this correctly first! :)

OpenStudy (anonymous):

Please! somebody :(

OpenStudy (anonymous):

a. The side opposite the 60 degree angle is given by the equation sin 60 = o / 41. This equation must be used since we are only give the length of one side. o is the number we're looking for and 41 is the given side which is adjacent to the 60 degree angle. The final solution to part a would then be (41x cos60) b. The hypotenuse can be found using the equation cos 60 = 41 / h. h of course stands for hypotenuse. Solving this equation for h gives us h = 41 / (cos60) c. This a pretty straight forward part. Simply compute sin30, cos30, and tan30. There a few ways to compute these, either using a table, a calculator, or a unit circle.

OpenStudy (anonymous):

Thanks! :)